# Restore Text First, Enhance Image Later: Two-Stage Scene Text Image Super-Resolution with Glyph Structure Guidance

### v1

@@ Abstract
Current generative super-resolution methods show strong performance on natural images but distort text, creating a fundamental trade-off between image quality and textual readability. To address this, we introduce TIGER (Text–Image Guided supEr-Resolution), a novel two-stage framework that breaks this trade-off through a “text-first, image-later” paradigm. TIGER explicitly decouples glyph restoration from image enhancement: it first reconstructs precise text structures and then uses them to guide subsequent full-image super-resolution. This glyph-to-image guidance ensures both high fidelity and visual consistency. To support comprehensive training and evaluation, we also contribute the UltraZoom-ST (UltraZoom-Scene Text), the first scene text dataset with extreme zoom ($\times$14.29). Extensive experiments show that TIGER achieves state-of-the-art performance, enhancing readability while preserving overall image quality.
Figure 1: We present TIGER (Text–Image Guided supEr-Resolution), a novel framework for scene text super-resolution. Its ‘text-first, image-later’ paradigm ensures accurate glyph restoration and consistently high overall image fidelity and visual quality. See more on the project page.
@@ 1 Introduction
Scene Text Image Super-Resolution is a critical problem in computer vision for its vital role in car navigation (Li & Cui, 2025), scene understanding (Kil et al., 2023; Deshmukh et al., 2024), and document enhancement (Souibgui et al., 2023). It seeks to restore a high-quality super-resolution (SR) image from a degraded low-resolution (LR) input while preserving the correct text meaning. Unlike general image super-resolution for natural scenes, where plausible texture synthesis or detail hallucination is acceptable and does not alter the semantic meaning of the image, visual texts exhibit extremely low tolerance to structural errors. This is especially evident in Chinese, where a minor stroke distortion or omission may entirely alter the meaning (Yu et al., 2023; Li et al., 2023a).
In recent years, with the rapid development of generative models such as diffusion (Ho et al., 2020; Dhariwal & Nichol, 2021; Saharia et al., 2022; Ramesh et al., 2022; Rombach et al., 2022; Chang et al., 2023; Zhang et al., 2024b), image super-resolution has increasingly relied on their strong generative priors to recover missing details from low-resolution (LR) inputs (Wu et al., 2024b; Yu et al., 2024; Lin et al., 2024; Wu et al., 2024a; Dong et al., 2025a; Duan et al., 2025; Hu et al., 2025). Although these approaches can effectively restore fine-grained natural textures (e.g., grass, leaves), they often distort text regions, turning them into gibberish as shown in Fig.1 (b). Some researchers attempt to address this by processing only text regions (Zhang et al., 2024a; Li et al., 2023a). While these methods improve readability notably, the absence of global background constraints often introduces style inconsistencies and block artifacts between text and background. The unsatisfactory performance of the present methods can be attributed to several reasons. Firstly, there is a deficiency in scene text paired data. Current scene text super-resolution datasets (Wang et al., 2020; Yu et al., 2021; Ma et al., 2023) are limited in degradation and focused on textline annotations, making it difficult for models to learn the mapping from low- to high-resolution text. Moreover, current methods primarily enhance overall image quality but fail to capture fine-grained glyph structures. This limitation is amplified for Chinese characters, due to their complex glyph designs and relatively low saliency in images. Consequently, models tend to under-represent and collapse glyphs into oversimplified, averaged forms, producing overlapping or distorted characters. Ultimately, current approaches face a persistent trade-off between maintaining readability and ensuring high image quality. Our key observation is that these goals need not be antagonistic if text and non-text are explicitly handled differently. Text structures can be reconstructed with dedicated mechanisms and then used to guide full-image restoration, ensuring coherent style without artifacts.
@@
Table 1: Dataset comparison. ‘Non-Latin characters’ indicates inclusion of languages such as Chinese, ‘Multi-lines’ denotes images with multiple text lines, ‘Real-world degradation ($\times 4$)’ refers to captured LR with zoom beyond $\times 4$. TextZoom (Wang et al., 2020) lacks non-Latin text, CTR (Yu et al., 2021) lacks multi-line text, and Real-CE (Ma et al., 2023) has only mild degradation. UltraZoom-ST covers all, providing a more challenging benchmark.
Content TextZoom CTR Real-CE UltraZoom-ST (Ours)
Non-Latin characters
Multi-lines
@@
Moreover, to address the issue of data scarcity and enable comprehensive evaluation (Wang et al., 2020; Ma et al., 2023), we introduce the UltraZoom-ST (UltraZoom-Sence Text) benchmark dataset, the first scene text dataset that contains extreme zooming mode ($\times 14.29$), providing extra challenging scenarios for the field. Content differences are outlined in Table 1. It includes high-quality 5,036 LR–HR pairs captured at multiple focal lengths (ranging from 14mm to 200mm), with 49,675 text lines in total. Each pair comes with detailed annotations, including detection boxes and text transcripts, to support both training and evaluation. The dataset includes diverse scenarios such as shop signs, posters, and documents. It also covers varying lighting conditions, including daylight, indoor lighting, and nighttime, offering a challenging yet practical benchmark for the field. To ensure reliable alignment across extreme focal lengths, we adopt a coarse-to-fine cascade alignment strategy: coarse alignment is performed via global geometric transformation, and residual misalignments are corrected through local refinement. This process achieves accurate LR–HR alignment, thereby ensuring both the quality of the dataset and the reliability of our experiments. The main contributions are summarized as follows:
@@
    We propose TIGER, the first two-stage scene text super-resolution framework that introduces a novel ‘text-first, image-later’ paradigm to decouple glyph restoration from image enhancement, improving both readability and visual quality.
@@
    We introduce UltraZoom-ST, the first scene text benchmark with extreme zooming ($\times$14.29), offering well-aligned, richly annotated LR–HR pairs for comprehensive evaluation under challenging and diverse real-world conditions.
@@
    Extensive experiments on both Real-CE and UltraZoom-ST show the proposed method outperforms prior state-of-the-art models, particularly in preserving text structure fidelity.
@@ 2 Related Works
Real-World Image Super-Resolution.  Real-world image super-resolution (Real-SR) aims to reconstruct high-resolution (HR) images from low-resolution (LR) inputs degraded under uncontrolled real-world conditions. Early GAN-based methods, such as BSRGAN (Zhang et al., 2021) and Real-ESRGAN (Wang et al., 2021), synthesize degradations through random combinations of known distortions and use adversarial training for restoration. While they generate natural-looking images, their instability and insensitivity to fine details limit their ability to recover structural elements, especially text. Recent work introduces diffusion models into Real-SR (Yue et al., 2023; Wang et al., 2024b), improving perceptual quality but still struggling with complex structures like scene text. StableSR (Wang et al., 2024a) and DiffBIR (Lin et al., 2024) apply ControlNet (Zhang et al., 2023) to condition generation on LR inputs, while PASD (Yang et al., 2024) and SeeSR (Wu et al., 2024b) incorporate high-level semantics to enhance fidelity. SUPIR (Yu et al., 2024) scales training with large image-text pairs and introduces degradation-robust encoders. Other approaches, including OSEDiff (Wu et al., 2024a) and TSD-SR (Dong et al., 2025b), directly apply the diffusion process on LR images and distill models for one-step sampling. DiT-SR (Cheng et al., 2025), DreamClear (Ai et al., 2024), and DiT4SR (Duan et al., 2025) adopt diffusion transformers (DiT) for Real-SR. Despite enhancing perceptual fidelity, these methods often overlook accurate text structures. TADiSR (Hu et al., 2025) addresses Chinese scene text super-resolution by using Kolors as the base model and aggregating cross-attention maps for text structure supervision. However, constrained by the resolution of cross-attention, it struggles to restore small or severely degraded text.
Text Image Super-Resolution. Text image super-resolution (Text-SR) restores textual content from cropped images containing isolated words or text lines. Early methods (Dong et al., 2015) apply general SR architectures such as SRCNN (Dong et al., 2014) to enhance OCR performance on LR. TextSR (Wang et al., 2019) introduces GANs with text recognition loss, while PlugNet (Mou et al., 2020) and STT (Chen et al., 2021) jointly train SR and recognition modules for more discriminative features. TSRN (Wang et al., 2020) introduces the TextZoom dataset and incorporates an edge-aware module to preserve text details, while TATT (Ma et al., 2022) uses a global attention module to handle irregular text layouts. Recent methods leverage stronger generative priors. MARCONet (Li et al., 2023a) employs StyleGAN priors and a glyph structure codebook for realistic text reconstruction. DiffTSR (Zhang et al., 2024a) employs latent diffusion to separately denoise text and text-image components. Despite progress in text structure restoration, existing methods lack global background constraints, causing style inconsistencies and block artifacts between text and background.
@@
Figure 2: The framework of TIGER, which includes the Text Restoration stage (stage 1) and the Image Enhancement stage (stage 2). Stage 1 refines text regions to recover accurate glyph structures. Stage 2 uses these structures to guide full-image restoration for coherent text and background.
@@ 3.1 Architecture Overview
Current image super-resolution methods focus on enhancing overall image quality but often fail to accurately preserve glyph structures, leading to distorted text in super-resolved images, as shown in Fig. 1 (b). Conversely, text image super-resolution methods improve text readability but do not retain the global semantic information of the background, causing incoherence between text and non-text regions. To leverage the strengths of both approaches, we introduce the TIGER framework. As illustrated in Fig. 2, the TIGER framework is composed of 2 stages, the Text Restoration stage and the Text Restoration stage. In the text restoration stage, we extract the text regions from the LR input $x_{L}\in\mathbb{R}^{H\times W\times C}$ and feed them into the glyph structure restoration model to restore the text structure based on the text region of the LR input and the predicted text. We then reassemble the text structures to their original positions to obtain a text mask $\hat{x}_{m}\in\mathbb{R}^{H\times W\times C}$. In the image enhancement stage, the text mask and LR input are then processed by a ControlNet-like network to obtain the enhanced SR output $\hat{x}_{H}\in\mathbb{R}^{H\times W\times C}$. The following sections detail each stage.
@@ 3.2 Text Restoration stage
Text Restoration Pipeline. Existing methods like Hi-SAM (Ye et al., 2024) work well on clean high-resolution text but fail on the incomplete, distorted text in low-resolution images. To address this, we propose a region-level text restoration pipeline shown in Fig. 2 built on Rombach et al. (2022) to recover glyph structures $x_{m}\in\mathbb{R}^{H\times W\times C}$. An OCR detector first localizes text regions $\{x^{0}_{L},...,x^{N-1}_{L}\}$ and extracts their contents $\{y^{0},...,y^{N-1}\}$ as semantic conditions. Each region $\tilde{x}_{L}$ is encoded by VAE (Kingma & Welling, 2013) into $\tilde{z}_{L}\in\mathbb{R}^{h\times w\times c}$, concatenated with noise $z_{T}\in\mathbb{R}^{h\times w\times 2c}$, and iteratively denoised by a UNet $\epsilon_{\theta}$ into two branches: $z^{RGB}_{t-1}\in\mathbb{R}^{h\times w\times c}$ for appearance and $z^{m}_{t-1}\in\mathbb{R}^{h\times w\times c}$ for structure, which are merged as $z_{t-1}$. The text content $y$ is embedded into $c_{te}$ and fused into $z_{t}$ via cross-attention to guide structure recovery. After $T$ denoising steps, the mask branch output $z^{m}_{0}$ is decoded into $\tilde{x}_{m}$, and all restored regions are assembled into the final text mask $\hat{x}_{m}$. Focusing on real text reconstruction enables the model to capture glyph structures without non-text interference and reduces sensitivity to text saliency.
@@
To improve the quality of the text mask, we propose a segmentation-oriented loss. Let $\mathcal{\varepsilon}_{t}$ denote the noise predicted by the denoiser network $\epsilon_{\theta}$. Following  Ho et al. (2020), $z^{m}_{0}$ can be estimated by combining the time step $t$ with the noisy latent image $z^{m}_{t}$. This estimate is subsequently passed through the VAE decoder to obtain an approximate reconstruction of the original input text mask, denoted as $x^{\prime m}_{0}$. In this way, text mask generation can be supervised at the pixel level. We combine Mean Squared Error (MSE), Focal, and Dice Losses, commonly used in segmentation tasks (Ye et al., 2024), to compare $x^{\prime m}_{0}$ with the original image $x^{m}_{0}$, as shown in the following equation:
$\mathcal{L}_{Seg}=\|x^{\prime m}_{0}-x^{m}_{0}\|^{2}_{2}+\lambda_{Focal}\text{FocalLoss}(x^{\prime m}_{0},x^{m}_{0})+\lambda_{Dice}\text{DiceLoss}(x^{\prime m}_{0},x^{m}_{0}),$ (3)
@@ 3.3 Image Enhancement stage
Image Enhancement pipeline. To effectively take advantage of the generated glyph structure $\hat{x}_{m}$ and enhance the quality of LR $x_{L}$, we adopt a ControlNet (Zhang et al., 2023) network $\epsilon_{\phi}$. As illustrated in Fig. 2, after getting their latent representation $\hat{z}_{m}$ and $z_{L}$, we use the network to denoise the $z_{L}$ at $t$ timestep, using the null-text embedding $c_{Null}$. The output super-resolved latent representation $\hat{z}_{H}$ will be denoted as:
@@
where $\sigma_{t}$ is a scalar determined by the predefined diffusion time step. Together with the text restoration stage, this image enhancement pipeline forms a progressive ‘text-first, image after’ paradigm; our design injects structure-aware control into the generative process, allowing the network to enhance global image quality without eroding the recovered text structures.
@@
where $\lambda_{l2}$ and $\lambda_{LPIPS}$ are balancing coefficients for different loss terms. To enhance glyph control and emphasize the glyph structure, we extract the text boundaries using Sobel operators (Roberts & Mullis, 1987). This edge loss is expressed as:
@@
Figure 3: Overview of UltraZoom-ST. (a) Real-CE LRs show only mild degradation (red box), while UltraZoom-ST LRs exhibit stronger degradation (red box), enabling a more comprehensive evaluation. (b) Coarse-to-fine alignment: images are sorted by focal length, each warped to the next higher-focal neighbor using an estimated homography matrix, then refined to the 200 mm GT.
Table 2: Evaluation results on image quality. Numbers in bold indicate the best performance, and underscored numbers indicate the second best. TIGER (Ours) performs best on image quality.
Methods Real-CE UltraZoom-ST

### v2

@@ Abstract
Current image super-resolution methods show strong performance on natural images but distort text, creating a fundamental trade-off between image quality and textual readability. To address this, we introduce TIGER (Text–Image Guided supEr-Resolution), a novel two-stage framework that breaks this trade-off through a “text-first, image-later” paradigm. TIGER explicitly decouples glyph restoration from image enhancement: it first reconstructs precise text structures and uses them to guide full-image super-resolution. This ensures high fidelity and readability. To support comprehensive training and evaluation, we present the UZ-ST (UltraZoom-Scene Text) dataset, the first Chinese scene text dataset with extreme zoom. Extensive experiments show TIGER achieves state-of-the-art performance, enhancing readability and image quality.
Figure 1: We present TIGER (Text–Image Guided supEr-Resolution), a novel framework for scene text super-resolution. Its ‘text-first, image-later’ paradigm ensures accurate glyph restoration and consistently high overall image fidelity and visual quality.
11footnotetext: Equal contribution.22footnotetext: Corresponding author.
@@ 1 Introduction
Scene Text Image Super-Resolution is a critical problem in computer vision for its vital role in car navigation [DBLP:conf/iclr/LiC25], scene understanding [kil2023prestu, deshmukh2024textual], and document enhancement [souibgui2023text]. It seeks to restore a high-quality super-resolution (SR) image from a degraded low-resolution (LR) input while preserving the text meaning. Unlike general image super-resolution for natural scenes, where plausible texture synthesis or detail hallucination is acceptable and does not alter the semantic meaning of the image, visual texts exhibit extremely low tolerance to structural errors. This is especially evident in Chinese, where a minor stroke distortion or omission may entirely alter the meaning [yu2023chinese, li2023learning].
In recent years, with the rapid development of generative models such as diffusion [ho2020denoising, dhariwal2021diffusion, saharia2022photorealistic, ramesh2022hierarchical, rombach2022high, chang2023muse, zhang2024artbank], image super-resolution has increasingly relied on their strong generative priors to recover missing details from low-resolution (LR) inputs [wu2024seesr, yu2024scaling, lin2024diffbir, wu2025one, DBLP:conf/cvpr/DongFGWZCLZ25, duan2025dit4sr, hu2025text]. Although these approaches can effectively restore fine-grained natural textures (e.g., grass, leaves), they often distort text regions, turning them into gibberish as shown in Fig.1 (b). Because their primary goal is overall image enhancement, they tend to overlook text structure, resulting in a limited ability to preserve fine-grained glyph details. This limitation is amplified for Chinese characters, due to their complex and diverse glyph designs and relatively low saliency in images. Consequently, models tend to under-represent and collapse glyphs into oversimplified, averaged forms, producing overlapping or distorted characters. Some researchers attempt to address this by processing only text regions [zhang2024diffusion, li2023learning]. While these methods improve readability notably, the absence of global background constraints often introduces style inconsistencies and block artifacts between text and background. Ultimately, current approaches face a persistent trade-off between maintaining readability and ensuring high image quality. Our key observation is that these goals need not be mutually exclusive if text and non-text are explicitly handled differently. Text structures can be reconstructed with dedicated mechanisms and then used to guide full-image restoration, ensuring coherent style without artifacts.
@@
Table 1: Dataset comparison. TextZoom [wang2020scene] lacks Chinese, CTR [yu2021benchmarking] lacks multi-line text, and Real-CE [ma2023benchmark] has mild degradation. UZ-ST covers all, enabling comprehensive evaluation.
Content TextZoom CTR Real-CE UZ-ST (Ours)
Chinese characters
Multi-line text
@@
Moreover, existing scene text super-resolution datasets [wang2020scene, yu2021benchmarking, ma2023benchmark] are limited to mild degradation ($\leq\!\times 4$), which fails to reflect real-world long-distance scenarios and restricts models from learning robust low- to high-resolution mappings. To address the issue of data scarcity and enable comprehensive evaluation [wang2020scene, ma2023benchmark], we introduce the UZ-ST (UltraZoom-Sence Text) benchmark dataset, the first scene text dataset that contains extreme zooming mode ($\times 14.29$), providing extra challenging scenarios for the field. Content differences are outlined in Tab. 1. It includes high-quality 5,036 LR–HR pairs captured at multiple focal lengths (ranging from 14mm to 200mm), with 49,675 text lines in total. Each pair comes with detailed annotations, including detection boxes and text transcripts, to support both training and evaluation. The dataset includes diverse scenarios such as shop signs, posters, and documents. It also covers varying lighting conditions, including daylight, indoor lighting, and nighttime, offering a challenging yet practical benchmark for the field. To ensure reliable alignment across extreme focal lengths, we adopt a coarse-to-fine cascade alignment strategy combined with manual filtering: coarse alignment is performed via global geometric transformation, and residual misalignments are corrected through local refinement. This process achieves accurate LR–HR alignment, thereby ensuring both the quality of the dataset and the reliability of our experiments. The main contributions are summarized as follows:
@@
    We propose TIGER, the first two-stage scene text super-resolution framework with a ‘text-first, image-later’ paradigm that decouples glyph restoration from image enhancement, improving readability and visual quality.
@@
    We introduce UZ-ST, the first scene text benchmark with extreme zoom ($\times$14.29), offering aligned, richly annotated LR–HR pairs for thorough evaluation under challenging real-world conditions.
@@
    Extensive experiments on Real-CE and UZ-ST show our method outperforms prior state-of-the-art models, especially in preserving text structure.
@@ 2 Related Works
Figure 2: The framework of TIGER, which includes the Text Restoration stage (stage 1) and the Image Enhancement stage (stage 2). Stage 1 recover accurate glyph structures from text regions. Stage 2 uses them to guide full-image restoration for coherent text and background.
Real-World Image Super-Resolution.  Real-world image super-resolution (Real-SR) aims to reconstruct high-resolution (HR) images from low-resolution (LR) inputs degraded under uncontrolled real-world conditions. Early GAN-based methods, such as BSRGAN [zhang2021designing] and Real-ESRGAN [wang2021real], synthesize degradations through random combinations of known distortions and use adversarial training for restoration. While they generate natural-looking images, their instability and insensitivity to fine details limit their ability to recover structural elements, especially text. Recent work introduces diffusion models into Real-SR [yue2023resshift, wang2024sinsr], improving perceptual quality but still struggling with complex structures like scene text. StableSR [wang2024exploiting] and DiffBIR [lin2024diffbir] apply ControlNet [zhang2023adding] to condition generation on LR inputs, while PASD [yang2024pixel] and SeeSR [wu2024seesr] incorporate high-level semantics to enhance fidelity. SUPIR [yu2024scaling] scales training with large image-text pairs and introduces degradation-robust encoders. Other approaches, including OSEDiff [wu2025one] and TSD-SR [dong2025tsd], directly apply the diffusion process on LR images and distill models for one-step sampling. DiT-SR [cheng2025effective], DreamClear [ai2024dreamclear], and DiT4SR [duan2025dit4sr] adopt diffusion transformers (DiT) for Real-SR. Despite enhancing perceptual fidelity, these methods often overlook accurate text structures. TADiSR [hu2025text] addresses Chinese scene text super-resolution by using Kolors as the base model and aggregating cross-attention maps for text structure supervision. However, constrained by the resolution of cross-attention, it struggles to restore small or severely degraded text.
Text Image Super-Resolution. Text image super-resolution (Text-SR) restores textual content from cropped images of isolated words or text lines. Early methods [dong2015boosting] apply general SR architectures such as SRCNN [dong2014learning] to enhance OCR performance. TextSR [wang2019textsr] introduces GANs with a recognition loss, while PlugNet [mou2020plugnet] and STT [chen2021scene] jointly train SR and recognition modules for more discriminative features. TSRN [wang2020scene] releases the TextZoom dataset and employs an edge-aware module to preserve text details, while TATT [ma2022text] uses global attention for irregular layouts. Methods like MTDM [wang2024multi] inpaint text but fail on complex scripts due to primitive architectures. MARCONet [li2023learning] employs StyleGAN priors and a glyph structure codebook for realistic text reconstruction. DiffTSR [zhang2024diffusion] employs latent diffusion to separately denoise text and text-image components. Despite progress in text structure restoration, existing methods lack global background constraints, causing style inconsistencies and block artifacts between text and background.
@@ 3.1 Architecture Overview
Current image super-resolution methods focus on enhancing overall image quality but often fail to accurately preserve glyph structures, leading to distorted text in super-resolved images, as shown in Fig. 1 (b). Conversely, text image super-resolution methods improve text readability but do not retain the global semantic information of the background, causing incoherence between text and non-text regions. To leverage the strengths of both approaches, we introduce the TIGER framework. As illustrated in Fig. 2, the TIGER framework is composed of 2 stages, the Text Restoration stage and the Image Enhancement stage. In the text restoration stage, we extract the text regions from the LR input $x_{L}\in\mathbb{R}^{H\times W\times C}$ and feed them into the glyph structure restoration model to restore the text structure based on the text region of the LR input and the predicted text. We then reassemble the text structures to their original positions to obtain a text mask $\hat{x}_{m}\in\mathbb{R}^{H\times W\times C}$. In the image enhancement stage, the text mask and LR input are then processed by a ControlNet-like network to obtain the enhanced SR output $\hat{x}_{H}\in\mathbb{R}^{H\times W\times C}$. The following sections detail each stage.
@@ 3.2 Text Restoration stage
Figure 3: Overview of UZ-ST (UltraZoom-Scene Text). (a) Real-CE LRs show only mild degradation (red box), while UZ-ST LRs exhibit stronger degradation (red box), enabling a more comprehensive evaluation. (b) Coarse-to-fine alignment: images are sorted by focal length, each warped to the next higher-focal neighbor using an estimated homography matrix, then refined to the 200 mm GT.
Text Restoration Pipeline. Existing methods like Hi-SAM [ye2024hi] work well on clean high-resolution text but fail on the incomplete, distorted text in low-resolution images. To address this, we propose a region-level text restoration pipeline shown in Fig. 2 built on [rombach2022high] to recover glyph structures $x_{m}\in\mathbb{R}^{H\times W\times C}$. An OCR detector first localizes text regions $\{x^{0}_{L},...,x^{N-1}_{L}\}$ and extracts their contents $\{y^{0},...,y^{N-1}\}$ as semantic conditions. Each region $\tilde{x}_{L}$ is encoded by VAE [kingma2013auto] into $\tilde{z}_{L}\in\mathbb{R}^{h\times w\times c}$, concatenated with noise $z_{T}\in\mathbb{R}^{h\times w\times 2c}$, and iteratively denoised by a UNet $\epsilon_{\theta}$ into two branches: $z^{RGB}_{t-1}\in\mathbb{R}^{h\times w\times c}$ for appearance and $z^{m}_{t-1}\in\mathbb{R}^{h\times w\times c}$ for structure, which are merged as $z_{t-1}$. The text content $y$ is embedded into $c_{te}$ and fused into $z_{t}$ via cross-attention to guide structure recovery. After $T$ denoising steps, the mask branch output $z^{m}_{0}$ is decoded into $\tilde{x}_{m}$, and all restored regions are assembled into the final text mask $\hat{x}_{m}$. Focusing on real text reconstruction enables the model to capture glyph structures without non-text interference and reduces sensitivity to text saliency.
@@
To improve the quality of the text mask, we propose a segmentation-oriented loss. Let $\mathcal{\varepsilon}_{t}$ denote the noise predicted by the denoiser network $\epsilon_{\theta}$. Following  [ho2020denoising], $z^{m}_{0}$ can be estimated by combining the time step $t$ with the noisy latent image $z^{m}_{t}$. This estimate is subsequently passed through the VAE decoder to obtain an approximate reconstruction of the original input text mask, denoted as $x^{\prime m}_{0}$. In this way, text mask generation can be supervised at the pixel level. We combine Mean Squared Error (MSE), Focal, and Dice Losses, commonly used in segmentation tasks [ye2024hi], to compare $x^{\prime m}_{0}$ with the original image $x^{m}_{0}$:
$\mathcal{L}_{Seg}=\|x^{\prime m}_{0}-x^{m}_{0}\|^{2}_{2}+\lambda_{Focal}\text{FocalLoss}(x^{\prime m}_{0},x^{m}_{0})$ (3)
$+\lambda_{Dice}\text{DiceLoss}(x^{\prime m}_{0},x^{m}_{0}),$
@@ 3.3 Image Enhancement stage
Image Enhancement pipeline. To effectively take advantage of the generated glyph structure $\hat{x}_{m}$ and enhance the quality of LR $x_{L}$, we adopt a ControlNet [zhang2023adding] network $\epsilon_{\phi}$. As illustrated in Fig. 2, after getting their latent representation $\hat{z}_{m}$ and $z_{L}$, we use the network to denoise the $z_{L}$ only at the specific timestep $t$, using the null-text embedding $c_{Null}$. The output super-resolved latent representation $\hat{z}_{H}$ will be denoted as:
@@
where $\sigma_{t}$ is a scalar determined by the predefined diffusion time step. With the text restoration stage, this image enhancement pipeline forms a progressive ‘text-first, image after’ paradigm. Our design injects structure-aware control into the generative process, allowing the network to enhance global image quality while preserving text structures.
@@
where $\lambda_{l2}$ and $\lambda_{LPIPS}$ are balancing coefficients for different loss terms. To enhance glyph control and emphasize the glyph structure, we extract the text boundaries using Sobel operators [roberts1987digital]. This edge loss is expressed as:
@@
4 Dataset and Benchmark
Existing datasets like ma2023benchmark offer only subtle degradation, as shown in Fig. 3 (a), making them insufficient for evaluating model robustness. To address the lack of both challenging and publicly available datasets tailored for scene text image super-resolution, especially for Chinese text, we introduce UZ-ST, a challenging real-world benchmark. It is collected using ViVO X200 Ultra equipped with four fixed focal lengths (14 mm, 35 mm, 85 mm, 200 mm), enabling image pairs with extreme $\times 14.29$ zoom. It includes diverse scenes—street views, book covers, advertisements, menus, and posters—captured under varied lighting, with all text lines manually annotated. However, such extreme zoom introduces severe misalignment that breaks pixel-wise optimization methods like [cai2019toward], originally designed for moderate $\times 2$ to $\times 4$ zoom. To address this, we design a Cascade Coarse-to-fine Alignment pipeline (see Fig. 3 (b)): images are first sorted by focal length, and each low-quality image is sequentially aligned to its next higher-focal neighbor using optimization-based registration or feature-based extractors [lowe1999object, ripe2025], then refined against the 200 mm ground truth. Finally, we manually filter or realign problematic cases to ensure precise alignment.
Table 2: Evaluation results on image quality and text accuracy. Numbers in bold indicate the best performance. TIGER performs best.
Methods Real-CE UZ-ST